3 pts] At a locus with three alleles in a population of voles, you accumulate the following genotypic data:

genotype

observed # individuals

A1A1

19

A2A2

11

A3A3

14

A1A2

56

A1A3

79

A2A3

51

total

230

Is this population in Hardy Weinberg equilibrium at this locus? Determine using a chi-square test.
Answer #1

Answer:

A1A1-19 = 19*2=38 A1 alleles

A2A2-11 = 22 A2 alleles

A3A3-14 = 28 A3 alleles

A1A2-56 = 56A1 & 56 A2 alleles

A1A3-79= 79A1 & 79 A3

A2A3-51 = 51 A2 & 51 A3 alleles

Total alleles = 460

Total A1 alleles = 38+56+79 = 173

Frequency of A1 allele = 173/460 = 0.376

Total A2 alleles = 22+56+51 = 129

Frequency of A2 allele = 129/460 = 0.28

Total A3 alleles = 28+79+51 = 158

Frequency of A3 allele = 158/460 = 0.344

Expected number of Frequency of each genotype:

A1A1 =0.376 * 0.376 * 230 = 32.52

A2A2 = 0.28 * 0.28 * 230 = 18.03

A3A3 = 0.344 * 0.344 * 230 = 27.22

A1A2 = 2* 0.376 * 0.28 *230 = 48.43

A1A3 = 2 * 0.376 * 0.344 *230 = 50.49

A2A3 = 2 * 0.28 * 0.344 *230 = 44.31

Phenotype

Observed(O)

Expected (E)

O-E

(O-E)2

(O-E)2/E

A1A1

19

32.52

-13.52

182.79

5.621

A2A2

11

18.03

-7.03

49.42

2.741

A3A3

14

27.22

-13.22

174.77

6.421

A1A2

56

48.43

7.57

57.30

1.183

A1A3

79

59.49

19.51

380.64

6.398

A2A3

51

44.31

6.69

44.76

1.010

Total

230

230.00

0

23.374

Chi-square value = 23.374

Degrees of freedom = number of phenotypes – 1

Df = 6-1=5

P value = 11.07

The X^2 value of 23.374 is greater than the critical (p) value of 11.07. We can reject null hypothesis and the population is not in H-W equilibrium